OpenStudy (anonymous):

Find the sum twelve and three-eighths plus two and one-fifth. Be sure to reduce to simplest form. fourteen and four-thirteenths fourteen and twenty-three fortieths twelve and twenty-three fortieths fourteen and four-fifths

OpenStudy (anonymous):

You need to first convert each mixed number to a fraction, so for example 12 and 3/8=99/8. Then you need to find the least common factor of the two denominators so you can add them together. So (99/8 + 11/5) = (495/40 + 88/40) = 583/8 = 14 23/40

OpenStudy (anonymous):

14 and 23/40*

OpenStudy (anonymous):

can u help me wth more

OpenStudy (anonymous):

You won't learn from it if I do them all for you. If you must finish them now, you can use this website to calculate them for you. http://www.calculatorsoup.com/calculators/math/mixednumbers.php

OpenStudy (anonymous):

k